OpenStudy (megannicole51):

Find the sum: 5-(5/3)+(5/9)-(5/27)+(5/81)-......

OpenStudy (megannicole51):

so I know the formula to find the ratio is Sn+1/Sn and i thought i found the ratio and i got -1/3 which is probably wrong :(

OpenStudy (wwe123):

To find out if a sequence is arithmetic when given 4 terms: 1. Subtract 2nd term minus 1st term. 2. Subtract 3rd term minus 2nd term. 3. Subtract 4th term minus 3rd term. If all three are the same then the sequence is arithmetic, with d = common difference = this value. To find out if a sequence is geometric when given 4 terms: 1. Divide 2nd term by 1st term. 2. Divide 3rd term by 2nd term. 3. Divide 4th term by 3rd term. If all three are the same then the sequence is geometric, with r = common ratio = this value. ----------------------- First we test to see if it's an arithmetic sequence: 1. Subtract 2nd term minus 1st term. (5/9) - (-5/3) = 20/9 2. Subtract 3rd term minus 2nd term. (-5/27) - (5/9) = -20/27 3. No need to subtract 4th term minus 3rd term since those aren't the same. We know it is not an arithmetic sequence. So that rules out choices 2 and 4. So it's between 1 and 3. Next we test to see if it's a geometric sequence: 1. Divide 2nd term by 1st term. (5/9)÷(-5/3) = (5/9)(-3/5) = -1/3 2. Divide 3rd term by 2nd term. (-5/27)÷(5/9)= (-5/27)(9/5) = -1/3 3. Divide 4th term by 3rd term. (5/81)÷(-5/27) = (5/81)(-27/5) = -1/3 All three are the same, so the sequence is geometric, with r = -1/3. The correct choice is 1.

OpenStudy (megannicole51):

so the sum is -1/3?

OpenStudy (megannicole51):

@wwe123

OpenStudy (wwe123):

yes

OpenStudy (megannicole51):

my online homework says its wrong:(

OpenStudy (megannicole51):

@DukeShadows

OpenStudy (anonymous):

i followed it all and it looked right to me...gosh im sorry

OpenStudy (megannicole51):

so 1/3 is the right answer? or do i need to do something with 1/3?

OpenStudy (megannicole51):

i thought i had to do more than just 1/3

OpenStudy (anonymous):

assuming thats the correct sequence then itd be right unless you had to plug it in as a function in that case

OpenStudy (anonymous):

@thomaster any ideas?

OpenStudy (megannicole51):

@agent0smith

zepdrix (zepdrix):

Hmmm I'm coming up with -5/4 for the answer. Can you check and see if that's right? I can try to explain it if it is. I'm not great with arithmetic and geometric series though :p

OpenStudy (megannicole51):

its not right:(

zepdrix (zepdrix):

Hmmm :c

ganeshie8 (ganeshie8):

since |r| < 1, we can use infinite sum formula i think

zepdrix (zepdrix):

bahhh i typed it into wolfram wrong!! :3 15/4? I cheats lol

OpenStudy (megannicole51):

yay!!!!!!! how did u do it????

zepdrix (zepdrix):

I can't remember how to find the common ratio and all that business... Refer to wwe's notes for that. I was simply looking at the sequence and tried to write it as a summation. See how it's alternating? So we'll have a (-1)^n somewhere in the summation giving us that alternating sign. And we have powers of 3 in the denominators (starting with 3^0 = 1 for the first term). \[\Large =\sum_{n=0}^{\infty}(-1)^n\frac{5}{3^n}\]Then from there I cheat and put it into wolfram :c http://www.wolframalpha.com/input/?i=summation+from+n%3D0+to+n%3Dinfinity+of+%28-1%29%5En+5%2F%283%5En%29

OpenStudy (megannicole51):

yeah thats the formula i used!! well i probably just worked it out wrong! and how did u put that in wolfram?!?

zepdrix (zepdrix):

Click the link, you can see the text I typed into the search box. You have to use a lot of WORDS to get Wolfram to work correctly.

OpenStudy (megannicole51):

can you type this into wolfram for me:))) hahaha